OpenStudy (anonymous):

i dnt think anyones gonna help but...hey if anyone can! can u help?!?! please i need this today right now lol:) at the begining of 20-month period, Stacie owns one clothing store. during that period,she opens a second clothing store in a deffernt location. the table shows the total monthly sales of stacies clothing stores for the 20-month period monthly sales(in thousands of dollars)[month 2 4 6 8 10 12 14 16 18 20] [sales 3 5 4 6 5 12 16 22 26 32] B. what are two equations that model the data? what are the domain and range of each equation?explain your process c. about how much money did stacies stores earn in the fifth month? how much money do u expect the stores to earn in the twenty-fourth month? explain

OpenStudy (anonymous):

could anyone please help need this today dont know what time all be going to bed please help

OpenStudy (anonymous):

what class is this for? grade level?

OpenStudy (anonymous):

This is a linear regression problem. You can fit two lines two the data, the first line for the points before she opened her second store, and the second line for the points after she opened her second store.

OpenStudy (anonymous):

are you supposed to do som kind of regression?

OpenStudy (anonymous):

u can model the trend of the real-world data in a scatter plot with an equation of a line. u can use the equationn to estimate or to make predictions.

OpenStudy (anonymous):

ok... so like @agdgdgdgwngo says, a linear regression.

OpenStudy (apoorvk):

I guess, Natashka, as you just said, plotting the scatter plot may help, atleast in imaging this thing.

OpenStudy (anonymous):

have you done linear regression before?

OpenStudy (anonymous):

if not, i guess your best bet is to plot these and make a best fit line by eyeballing it...

OpenStudy (anonymous):

Or you can do this entire question in a TI graphing calculator.

OpenStudy (anonymous):

or try this.. http://www.alcula.com/calculators/statistics/linear-regression/

OpenStudy (kropot72):

The average two-monthly sales for the first 10 months =\[\frac{1}{5}(3+5+4+6+5)=4.6\] When the two-monthly sales for the second ten months are plotted against time a nearly linear graph results with a positive slope of 4 per two month period. Therefore an equation of the form y = mx + b can be formed and the sales projected to 24 months.

OpenStudy (anonymous):

THANKS EVERYONE SORRY FOR NOT REPLYING I WAS WORKING ON IT GOOD NIGHT

ganeshie8 (ganeshie8):

natashka once you finish compare your equations against mine : if you need explanation i will send you. this is pretty involved B. equation1 : y = .25x + 3 doman : [2, 10] range : [3.5, 5.5] equation2 : y = 2.5x - 18 domain : [12, 20] range : [12, 32] C. using above equations, 5th month earnings : 0.25 x 5 + 3 = 4.25 24th month forecast : 2.5 x 24 - 18 = 42